OpenStudy (tiffanymak1996):

how to find the interval of convergence of tan^-1 x

OpenStudy (tiffanymak1996):

the maclaurin series of tan^-1 x is: \[x-\frac{ x ^{3} }{ 3! }+\frac{ x^5 }{5}-\frac{ x^7 }{ 7 }+...\]

OpenStudy (anonymous):

I assume you mean the interval of convergence for the power series representation, right?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

Okay, so in general, we have that the terms of the series are \[a_n = \frac{x^{2n+1}}{2n+1} \] do you agree?

OpenStudy (tiffanymak1996):

yes

OpenStudy (tiffanymak1996):

wait, what about the -1^k?

OpenStudy (anonymous):

Oops, add that in there, my mistake.

OpenStudy (anonymous):

To find the interval of convergence, we will demand that \[\lim_{n \rightarrow \infty} \left| \frac{a_{n+1}}{a_n}\right| < 1 \] When we plug in, we get: \[ \lim_{n\rightarrow \infty} x^2 \left( \frac{2n+1}{2n+3} \right) = x^2 < 1 \] so the power series is valid for all x such that -1 < x < 1.

OpenStudy (tiffanymak1996):

where hasthe (2n+1)/(2n+3) gone?

OpenStudy (tiffanymak1996):

?

OpenStudy (experimentx):

\[ a_n = \frac{x^{2n+1}}{(2n+1)!}\] \[ {a_{n+1} \over a_n} = \frac{x^{2n+3}}{(2n+3)!}\times \frac{(2n+1)!}{x^{2n+1}} = {x^2\over (2n+2)(2n+3)} \] converges for all values of x.

OpenStudy (tiffanymak1996):

ok, thx!

OpenStudy (anonymous):

No, that's not true. There's no factorial.

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@tiffanymak1996 The series converges for |x| < 1 . The fraction you asked about disappears because as you take the limit as n -> infinity, that fraction becomes 1.

OpenStudy (experimentx):

hold on ... i;ven't check looks like you are right http://www.wolframalpha.com/input/?i=expand+tan^-1+x+at+x%3D0

OpenStudy (experimentx):

for x=1